# Solving Oscillator ODEs via Soft-constrained Physics-informed Neural Network with Small Data

### v1

@@ 2.2.3 Energy Conservation Regularization Improves Result of Solving Duffing Oscillator
Where, $t\in[0,1.5]$, take $\alpha=1.0,\beta=1.0$ with initial condition $u_{0}=15,u^{\prime}_{0}=0$. Again, the exact solution is obtained by Runge-Kutta method. The energy conservation term of Eq. (8) is $E=\frac{1}{2}(\frac{du}{dt})^{2}+\frac{1}{2}\alpha u^{2}+\frac{1}{4}\beta u^{4}$, substituting to the regularization term $\mathcal{L}_{reg}$.
Considering the minimalistic data case, for second-order ODEs only two training points containing initial value are used, and the number of collocation points $N_{c}=40$ is chosen uniformly. The optimal tuning results of the PINN solving Duffing oscillator with and without the energy conservation regularization term are shown in Fig. 7(a) and (b), respectively. The PINN without energy conservation regularization in the minimalistic data case, which is affected by nonlinearity, fails to complete convergence throughout the entire 72000 epoches of training; In contrast, the PINN with energy conservation regularization achieves a significant improvement, which is in perfect agreement with the exact solution. It should be noted that energy conservation regularization is useful for Duffing because Duffing follows conservation of energy; It is useless for VDP because VDP is energy dissipative.
(a) w/o Energy Conservation Regularization
(b) with Energy Conservation Regularization
Figure 7: Results of PINN Solving Duffing Oscillator with Energy Conservation Regularization or Not.
@@ 3.1 Summary
(1) The success of deep neural networks relies on big data and strong computational power, yet neglects to fully utilize prior information or knowledge. In particular, it lacks robustness or even fails to converge in the case of small data. PINN is precisely a powerful complement to deep neural networks in small data scenarios. It inherits the strong expressive ability of NN and incorporates physical information and knowledge into neural network topology or learning architecture as well as computational process by means of hard-coding or soft constraints, which reduces data redundancy, improves generalization performance, and consequently increases computational efficiency, robustness, and interpretability. More specifically, PINN generalizes well even with only a small amount of (supervised or labeled) training data, remains robust to data imperfections (e.g., noise, missing or outliers), and provides accurate and physically consistent predictions that can even be used for extrapolation tasks.
@@
On the other hand, compared to numerical discretization methods (e.g., Runge-Kutta method, Finite Element Method), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods. Thus PINN is becoming a favorable catalyst for the emerging era of Digital Twins [ZZ23, YV23, YKH+24, QLWM24, SWvDP24].
@@
(2) We generalized the mathematical framework and computational flow of the soft-constrained PINN method for solving differential equations (e.g., ODEs/PDEs). It is clear from the composition of the loss function that PINNs are driven by both data and physical information. Compared to conventional NN, PINN adds differential equations (including initial, boundary conditions) and a priori information or knowledge (e.g., the law of energy conservation) as constrains or regularization. The working mechanism of the soft-constrained PINN method and its accuracy and efficiency were experimentally verified by solving typical (covering first and second order, linear and nonlinear) oscillator ODEs such as Primer, Van der Pol, and Duffing oscillators.
@@
PINN greatly reduces the need for labeled data. When the nonlinearity of the oscillator ODEs (e.g., Primer, Duffing) is weak, a very small amount of supervised training data plus a small amount of unsupervised collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (including initial values) are needed for first- or second-order ODEs, respectively. This is just as classical analytic or numerical integration methods require only one and two initial conditions, respectively, for solving first- and second-order differential equations. Strongly nonlinear oscillator ODEs (e.g., VDP) also require only an appropriate increase in the number of training and collocation points, which still has significant advantages over conventional NN.
with the aid of collocation points and the use of physical information or knowledge, PINN has the ability to extrapolate data outside the time domain covered by the training set, and is robust to noisy data, thus with enhanced generalization capabilities. The training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. Small data allows fewer network parameters, so it can be further accelerated.
@@
The second-order nonlinear Duffing oscillator regularized by energy conservation is able to converge quickly with minimalist data (2 training points containing initial values plus a few collocation points), whereas it does not converge without the energy conservation regularization. The case study results show that, the soft-constrained PINN method can easily impose a physical law (e.g., energy conservation) constraint by adding a regularization term to the total loss function, thus improving the solution performance of ODEs that obey this physical law.
@@
(3) The DeepXDE-based implementation of the PINN method is not only light code but also efficient in training. The acceleration effect of GPU on PINN is obvious, as PINN does not depend on big data, it has high cross-platform flexibility and can also be efficiently calculated on CPU platform.
@@ 3.2 Future Work
Based on this work, the next step could be to continue the exploration of PINN methods (both hard-coding and soft-constrained) for solving stiff ODE and PDE, as well as the exploration of sampling methods [WZT+23], which are critical to the accuracy and efficiency of the solution. In addition, PINN combined with domain decomposition can be extended to large problems [KKL+21], and further extended to solve integral differential equations (IDEs), fractional order differential equations (FDEs) and stochastic differential equations (SDEs) [LMMK21].
@@
The code will be available at https://github.com/mikelu-shanghai/PINNtoODEwithSmallData.

### v2

@@ 2.2.3 Conservation of Energy Regularization Improves Result of Solving Duffing Oscillator
Where, $t\in[0,1.5]$, take $\alpha=1.0,\beta=1.0$ with initial condition $u_{0}=15,u^{\prime}_{0}=0$. Again, the exact solution was obtained by Runge-Kutta method. The conservation of energy term of Eq. (8) is $E=\frac{1}{2}(\frac{du}{dt})^{2}+\frac{1}{2}\alpha u^{2}+\frac{1}{4}\beta u^{4}$, substituting to the regularization term $\mathcal{L}_{reg}$.
Considering the minimalistic data case, for second-order ODEs only two training points containing initial values were used, and the number of collocation points $N_{c}=40$ was chosen uniformly. The optimal tuning results of the PINN solving Duffing oscillator with and without the conservation of energy regularization term are shown in Fig. 7(a) and (b), respectively. The PINN without the conservation of energy regularization in the minimalistic data case, which is affected by nonlinearity, fails to complete convergence throughout the entire 72000 epoches of training; In contrast, the PINN with the conservation of energy regularization achieves a significant improvement, which is in perfect agreement with the exact solution. It should be noted that the conservation of energy regularization is useful for Duffing because Duffing follows the conservation of energy; It is useless for VDP because VDP is energy dissipative.
(a) w/o the Conservation of Energy Regularization
(b) with the Conservation of Energy Regularization
Figure 7: Results of PINN Solving Duffing Oscillator with the Conservation of Energy Regularization or Not.
@@ 3.1 Summary
(1) The success of deep neural networks relies on big data and strong computational power, yet neglects to fully utilize prior information or knowledge. In particular, it lacks robustness or even fails to converge in the case of small data. PINN is precisely a powerful complement to deep neural networks in small data scenarios. It inherits the strong expressive ability of neural networks and incorporates physical information and knowledge into neural network’s topology or learning architecture as well as computational process by means of hard-coding or soft constraints, which reduces data redundancy, improves generalization performance, and consequently increases computational efficiency, robustness, and interpretability. More specifically, PINN generalizes well even with only a small amount of (supervised or labeled) training data, remains robust to data imperfections (e.g., noise, missing or outliers), and provides accurate, fast and physically consistent predictions that can even be used for extrapolation tasks.
@@
On the other hand, compared to traditional numerical discretization methods (e.g., Runge-Kutta method, FEM), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods. Thus PINN is becoming a favorable catalyst for the emerging era of Digital Twins [ZZ23, YV23, YKH+24, QLWM24, SWvDP24].
@@
(2) We generalized the mathematical framework and computational flow of soft-constrained PINN for solving ODEs/PDEs. It is clear from the composition of the loss function that PINN is driven by both data and physical information. Compared to conventional NN, PINN adds differential equations (including initial, boundary conditions) and a priori information or knowledge (e.g., the law of conservation of energy) as constrains or regularization. The working mechanism of the soft-constrained PINN and its accuracy and efficiency were experimentally verified by solving typical (covering first- and second-order, linear and nonlinear) oscillator ODEs such as Primer, Van der Pol, and Duffing oscillators.
@@
PINN greatly reduces the need for labeled data. When the nonlinearity of the oscillator ODEs (e.g., Primer, Duffing) is weak, a very small amount of supervised training data plus a few unsupervised collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (including initial values) are needed for first- or second-order ODEs, respectively. This is just as classical analytic or numerical integration methods require only one and two initial conditions, respectively, for solving first- and second-order differential equations. Strongly nonlinear oscillator ODEs (e.g., VDP) also require only an appropriate increase in the number of training points, which still has significant advantages over conventional NN. With the aid of collocation points and the use of physical information or knowledge, PINN has the ability to extrapolate data outside the time domain of the training set, and is robust to noisy data, thus with enhanced generalization capabilities. The training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. Small data allows fewer network parameters, so it can be further accelerated.
@@
The second-order nonlinear Duffing oscillator regularized by conservation of energy is able to converge quickly with minimalist data (2 training points containing initial values plus a few collocation points), whereas it does not converge without the conservation of energy regularization. The case study results show that, soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law.
@@
(3) The DeepXDE-based implementation of PINN is not only light code but also efficient in training. The acceleration effect of GPU on PINN is obvious, as PINN does not depend on big data, it has high cross-platform flexibility and can also be efficiently calculated on CPU platform.
@@ 3.2 Future Work
Based on this work, the next step could be to continue the exploration of PINN methods (both hard-constrained and soft-constrained) for solving stiff ODEs and PDEs, as well as the exploration of sampling methods [WZT+23, YKH+24], which are critical to the accuracy and efficiency of the solution. In addition, PINN combined with domain decomposition can be extended to large problems [KKL+21], and further extended to solve integral differential equations (IDEs), fractional order differential equations (FDEs) and stochastic differential equations (SDEs) [LMMK21].
@@
The codes will be available at https://github.com/mikelu-shanghai/PINNtoODEwithSmallData. Readers in need can contact us.